OpenStudy (waterbottleisblue):

A circle is centered at (−4, −6) and has a radius of 3. Which of the following is the equation for this circle? a. (x + 4)2 + (y + 6)2 = 3 b. (x − 4)2 + (y − 6)2 = 9 c. (x − 4)2 + (y − 6)2 = 3 d. (x + 4)2 + (y + 6)2 = 9

OpenStudy (anonymous):

eq. of circle \[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] where (h,k) is the center and r be the radius of the circle.